OpenStudy (anonymous):

Evaluate the limit assuming that lim g(x)=1as x approaches -3: what is lim g(x)/x^2 as x approaches -3?

jimthompson5910 (jim_thompson5910):

\[\Large \lim_{x \to -3} g(x) = 1\]

jimthompson5910 (jim_thompson5910):

\[\Large \lim_{x\to -3}\left[\frac{g(x)}{x^2}\right] =\left[\frac{\displaystyle \lim_{x \to -3} g(x)}{\displaystyle \lim_{x \to -3} x^2}\right] \]

jimthompson5910 (jim_thompson5910):

I'll let you finish up

OpenStudy (anonymous):

i got 1/9 . thanks!

jimthompson5910 (jim_thompson5910):

correct